a rope is 21 yards long. marcos cut it into two pieces so that one piece is 3 yards longer than the other. find the length of each piece
Answer by jim_thompson5910(35256)

x+y = 21
y = x+3

--------------

x+y = 21

x+x+3 = 21

2x+3 = 21

2x = 21-3

2x = 18

x = 18/2

x = 9

y = x+3

y = 9+3

y = 12

-------------------------------------------------------

Final Answer:

The two pieces are 9 yards and 12 yards long.